Find the future value of an ordinary annuity if payments are made in the amount R and interest is compounded as given. Then determine how much of this value is from contributions and how much is from interest. R​;500 6.32​% interest compounded semiannually for 12 years.
Answer #1

Future value of annuity = PMT * (1 +R)^N - 1) / R

where PMT = 500, R = 6.32%/2 = 0.0316 and N = 12 *2 = 24

= 500 * ((1 + 0.0316)^24-1) / 0.0316

= 17562.49

Total contribution = 500 * 24 = 12000

Total interest = 17562.49 - 12000 = 5562.49